OpenStudy (anonymous):

How many different seven-digit numbers can be produced by rearranging the digits of 2502254 ?

OpenStudy (anonymous):

you have seven digits there, and so if they were all different, the answer would be \(7!\)

OpenStudy (anonymous):

but since you cannot tell one "2" from another, you have to divide that by \(3!\) your answer is therefore \[\frac{7!}{3!}\]

OpenStudy (anonymous):

but how do you know that none of those numbers are repeats?

OpenStudy (anonymous):

i thought you have to do it by making 0,2,4,5 each leading numbers and going from there

OpenStudy (anonymous):

since it said "rearranging" that is what i assumed

OpenStudy (anonymous):

oh damn my answer was wrong anyway!

OpenStudy (anonymous):

i think the answer is 360...i looked it up but it didn't make sense to me.

OpenStudy (anonymous):

you have two fives as well, so it is \[\frac{7!}{3!2!}\] do you understand this notation ?

OpenStudy (anonymous):

i get 420

OpenStudy (anonymous):

kind of....why is it only over 3! and 2! though?

OpenStudy (anonymous):

ok lets imagine we had 7 completely different number, say the digits 1,2,3,4,5,6,7 and we wanted to know how many ways we could arrange them. 7 choices for the first digit, then 6 for the second (since we used up one) then 5 for the third, then 4 then 3 then 2 then 1 by the "counting principle" the number of ways we can do these things all together is \[7\times 6\times 5\times 4\times 3\times 2\times 1 =7!\]

OpenStudy (anonymous):

so if they were all different, the number of ways to arrange them would be \(7!\) which is just shorthand for what i wrote above

OpenStudy (anonymous):

http://mvtrinh.wordpress.com/ i found this online and i have no idea what he did. i understand the way you're expaining it but i'm not sure which is right

OpenStudy (anonymous):

well this looks more complicated that what i wrote, but let me read it for a second

OpenStudy (anonymous):

okay thanks so much!

OpenStudy (anonymous):

aah ok lets continue with the method i wrote. the answer here is correct, but i think needlessly complicated

OpenStudy (anonymous):

going back to the counting principle, we see if they are all distinct there would be \(7!\) different arrangements. however, since we cannot tell the three "2"s apart, we have counted by \(3!=6\) too many ways. is that clear?

OpenStudy (anonymous):

that is the number of ways you can "permute" the three "2"s

OpenStudy (anonymous):

yeah i got that

OpenStudy (anonymous):

similarly, we cannot tell the two "5"s apart, so we have overcounted also by a factor of \(2!=2\)

OpenStudy (anonymous):

therefore i thought the answer was \[\frac{7!}{3!2!}=420\] but i made a MISTAKE! i forgot that you cannot begin a 7 digit number with zero!!

OpenStudy (anonymous):

ohhh! okay i get the zero part..but how would you account for that?

OpenStudy (anonymous):

so we subtract the ones we counted with a zero at the beginning

OpenStudy (anonymous):

how many are there? by the exact same reasoning, now that i have put zero in the first spot, there would be \(\frac{6!}{3!2!}=60\) possibilities, therefore we have over counted by 60

OpenStudy (anonymous):

and low an behold \(420-60=360\) just like it says in the book i totally ignored that zero business, but now we both learned something

OpenStudy (anonymous):

good catch by the way, finding that link on the other hand i think it is more complicated that necessary. on the other hand, as usual with math, there is a) more than one way to skin a cat and b) different interpretations depending on your point of view

OpenStudy (anonymous):

haha ! thank you soooooo much! you're a life saver!

OpenStudy (anonymous):

glad to help do you have to write this up and turn it in, or just get the answer, or none of the above (just curious)?

OpenStudy (anonymous):

write it up and turn it in

OpenStudy (anonymous):

ok, then don't lose this thread! should be clear now though, have fun

OpenStudy (anonymous):

i won't haah thanks again!

OpenStudy (anonymous):

yw (again)